# Modules as effective nodes in coarse-grained networks of Kuramoto oscillators

### v1

@@ III Two modules
Figure 2: Dots represents the numerical integration of system (4) for 3 different connection probabilities $p$ and frequency offsets $\omega$. (a): Critical inter-module coupling strength. Lines represent a 2 oscillators system. (b): Minimum inner-module coupling strength at phase transition. Lines simply connect the points. Global values: $N_{1}=200$ and $N_{2}=100$. Gaussian frequency distributions with $\Delta=1$ and $(\omega_{1},\omega_{2})=(1.5,0)$. Error bars made over 3 simulations.
IV Three modules
We now extend the results of the previous section to systems with three modules. We first derive an expression for $r$ when $N=3$. The Kuramoto dynamics can be written in terms of the phase differences $\phi_{12}\equiv\theta_{1}-\theta_{2}$ and $\phi_{23}\equiv\theta_{2}-\theta_{3}$ as
@@ IV Three modules
Similar to Eq. (8), the order parameter can be written in terms of phase differences as
@@
Figure 3 (a) shows the contour plot of Eq. (15) in the phase space $\phi_{12}\times\phi_{23}$. To determine the stability of these fixed points we calculated the Jacobian of system (13) and looked at the largest real part of the pair of eigenvalues: when it is zero we are at the critical line, shown as the red curve in Fig.3 (a) . The curve divides the phase space into two regions, inside where the system displays stable synchrony, and outside which presents periodic motion. More elucidating, however, is to understand synchronization in the parameter space $\Omega_{1}\times\Omega_{3}$. Figure 3 (b) shows the stable $r$ values from panel (a) reparameterized according to Eq. (14). It is very interesting to note that, regardless of being in the phase space or the parameter space, contours of constant synchronization resemble ellipses. We proceed to finding approximations for such contours. This will help fit simulations from three-module networks with the three-oscillator system.
Figure 3: (a): Eq. (15) in the phase space $\phi_{12}\times\phi_{23}$. Red curve divides stable from unstable points, stable within. (b): Stable points reparameterized to the parameter space $\Omega_{1}\times\Omega_{3}$ via Eq. (14). (c): How stable $r$ values changes through the ellipses. Dashed lines are the empirical values from Eq. (18). (d): Contrast between the ellipses (17) (dashed) and constant $r$ values (continuous).
@@
In the inset of Figure 3 (d) we can easily see that there are constant $r$ values that fall out of the theoretical ellipses, therefore, there should be a smaller ellipse that encompass all values. We numerically find that further increasing $\alpha$ up to $\alpha_{2}\equiv 9.295$ causes all possible synchronization outcomes to tightly fall within the ellipse. In other words, depending on how we control our parameters $\{\Omega_{1},\Omega_{2}\}$ the phase transition happens somewhere between $\alpha_{1}\leq\alpha\leq\alpha_{2}$, or in terms of the original parameters, $\lambda^{2}\leq\omega_{1}^{2}-\omega_{1}\omega_{3}+\omega_{3}^{2}\leq 1.0328\lambda^{2}$. Still, we have no approximation for $r$ in such regime.
@@
Figure 4: (a)-(c): Final $\lambda_{\text{in}}$, global order parameter $r$ and standard deviation $\sigma$ in the frequency parameter space. Simulation values: $N_{1}=90$, $N_{2}=100$, $N_{3}=110$, $p=0.01$ and $\lambda=3$. Gaussian frequency distributions with $\Delta=1$ and means $\omega_{1}$, $\omega_{2}=0$ and $\omega_{3}$.
@@
We now proceed to simulate Eq. (4) with three modules, as before we take modules of distinct sizes to be more realistic, namely $N_{1}=90$, $N_{2}=100$ and $N_{3}=110$. Throughout this section we fixed $p=0.01$ and $\lambda=3$ so that $\Omega_{i}=\omega_{i}$. Again, we set the frequency distributions to be Gaussians of width $\Delta=1$ and centers $\omega_{1}$, $\omega_{2}=0$ and $\omega_{3}$. The minium synchrony of each module is $q_{\sigma}=0.9$.
As before, we calculate the standard deviation of $r$. Figures 4 (b)-(c) respectively show the final value of the global order parameter and its standard deviation. It is also plotted in black the elliptical boundary of Eq. (17). We observe that the synchronization region corresponds to the expected behavior for three oscillators and exhibit all distinct behaviors, full and partial phase locking and asynchrony.
@@
For three modules we have seen so far that given high enough $\lambda_{\text{in}}$ the coarse-graining procedure retain the basic properties of the reduced system; now we discuss how different parameter values affect the internal coupling. Figure 4 (a) shows the final value of $\lambda_{\text{in}}$. We see the same behavior of two modules, approaching the synchronization regime from a asynchronous state increases the internal coupling required to preserve $r_{\sigma}\geq q_{\sigma}$. This might seem an odd behavior considering we are decreasing the absolute frequency value and consequently expect it to be easier to keep cohesive modules. However, given that there are few connections between modules, the transfer of information from one to the others is harder when their frequencies are further apart, even though the absolute value is smaller.
V Real Networks
@@
So far we have seen conditions for the known behavior of 2 and 3 oscillators emerge in artificial modular networks. These networks, however, were generated in such a way as to exhibit very strong modular structure that does not resemble real world systems. In what follows we test our results on two famous networks, Zachary’s Karate club social network and the C. Elegans gap junctions neural network [46, 26]. Both of these networks can be modularized. What distinguishes them from the artificial networks constructed in this paper are structural properties such as not having full number of connections within modules and non random connections between modules per se. We show the effects of the coarse-graining procedure in these networks.
@@
Figure 5: Top: Dynamics of Eq. (4) for Zachary’s Karate social network. (a): Final $\lambda_{\text{in}}$. Magenta and turquoise lines marks the average of critical $\lambda$ and $\lambda_{\text{in}}$. (b): $\langle r\rangle$ with $\lambda_{\text{in}}$ relative to (a) in purple and with fixed $\lambda_{\text{in}}=20$ in red, green ribbon represents Eq. (11) for $0.9\leq q_{\sigma}\leq 1$. Simulation values: $\Delta=0$, $\omega_{1}=1.5$ and $\omega_{2}=0.0$. Bottom: Dynamics of Eq. (4) for C. Elegans gap junctions network. (c)-(e): $\langle r\rangle$ for the network with fixed $\lambda_{\text{in}}=20$ in red, equivalent coarsed-grained system in green. Simulation values: $\Delta=0$, (a): $(\omega_{1},\omega_{2},\omega_{3})=(0,1,6)$; (b): $(\omega_{1},\omega_{2},\omega_{3},\omega_{4},\omega_{5})=(0,1,1,-2,7)$; (c): $(\omega_{1},\omega_{2},\omega_{3},\omega_{4},\omega_{5},\omega_{6},\omega_{7},\omega_{8},\omega_{9},\omega_{10})=(0,1,-3.5,-2,2,0.6,0.4,1.2,1.5,-1.5)$.
@@
Let us start with the Karate network [46]. It has a total of 34 nodes with 78 connections divided into two modules of 17 nodes each. As before, we put this network under the dynamics given by Eq. (4) and slowly increase $\lambda_{\text{in}}$ until a minimum inner module synchrony of 90% is reached. Figure 5 (a) shows the minimum internal coupling that keeps the internal synchrony threshold immediately after the phase transition. As expected, increasing $\lambda$ for asynchronized states impairs modular synchrony, hence the continuous increase in $\lambda_{\text{in}}$ as $\lambda$ grows. On the other hand, after the phase transition, synchronized states act in cohesion, requiring lesser $\lambda_{\text{in}}$ as we increase $\lambda$. Note that the phase transition is at the expected value of 1.5 and thus is not affected by the structural differences of Zachary’s network. However, when comparing Zachary’s network to the artificial network used in the 2 modules section, synchronized states do not reach the synchronization values of 2 oscillators system, this can be seen in Figure 5 (b) where the purple curve shows the $r$ values for Zachary’s network, note that they do not fall within the 2 oscillators ribbon shown in green. Therefore, the structural differences between the real and artificial networks are affecting the coarse-graining procedure. We can counteract the effects of network structure by overshooting the internal coupling, the red curve of Figure 5 (b) shows the synchronization values when keeping $\lambda_{\text{in}}=20$ throughout, this forces the system to behave as 2 oscillators in spite of networks effects.
As for our second example we study a reduced network from the C. Elegans gap junctions neural network containing 248 nodes and 511 connections. Using appropriate metrics it can be modularized into 3, 5 or 10 modules [47, 48]. Note that regardless of number of modules, some end up being far greater than others, for instance, in the 3 modules arrangement we have 130 nodes in the largest module, 77 nodes in the intermediate and 41 nodes in the smallest, for 5 modules the number of nodes in each module are 120, 83, 34, 7 and 4 and with 10 modules the range goes from 76 nodes to only 6. Add to that the fact that the network has about 1% of the total number of possible connections to conclude that it is very hard to make all modules have a minimum synchrony of 90% or even lower. For that reason we choose only to overshoot $\lambda_{\text{in}}$ and compare whether the global order parameter of the network resembles that of the coarse-grained system for the cases of 3, 5 and 10 oscillators. Figure 5 (c)-(e) shows in red the simulation of $\langle r\rangle$ for those 3 modularizations, each accompanied by a simulation of the order parameter of the reduced system in green for comparison. For 3 modules results show that despite the lack of oscillatory behavior by the neural network, exposed by the low standard deviation ribbon, synchronization values fall somewhat in the mean values of the coarse-grained system, the same is valid for 10 modules. Differently, 5 modules achieve global synchronization much faster in the neural network then in the reduced system, this is most likely due to the modularization process. As mentioned earlier, the C. Elegans gap junctions neural network is biased, not all modules are equally sized, however, specifically for 5 modules the discrepancy in amount of nodes in each group is much higher then in the other settings, this causes the system to be highly dependent on very specific modules which, if synchronized enough trough $\lambda_{\text{in}}$ values, already gives global synchronization regardless of other modules.
VI Conclusions
In this work we studied a coarse-graining process in modular networks where each module is reduced to a single effective node. We constructed artificial networks where modules are fully connected and nodes interact with strength $\lambda_{\text{in}}$ for intra-modular connections, and $\lambda$ for inter-module connections, that exist with probability $p$. Using Kuramoto oscillators as underlying dynamics we derived the Eq. (6) for the average dynamics of the modules, which is independent of the network details. This constitutes the base of our coarse-graining process, leading to the approximation where the dynamics of $s$ modules would be equivalent to that of a system composed of $s$ oscillators. We tested this prediction by simulating the Kuramoto model with 2 and 3 modules. Moreover, we included the assumption of high internal synchrony into the estimate of the order parameters via $r_{\sigma}\geq q_{\sigma}$. Thus, our coarse-graining procedure gives a lower bound for the global network behaviors when the system is replaced by $s$ oscillators, where $q_{\sigma}$ is the minimum modular synchrony.
Results of the coarse-graining method are shown in Figures 1 and 4 for 2 and 3 modules. Figures 1(b) and 4(a) show that, starting from global asynchrony but keeping $r_{\sigma}\geq q_{\sigma}$, approaching the first order phase transition demands increasingly larger inner-module coupling strength. In this regime the connections among the modules, responsible for bringing them together, hinder internal module’s synchrony. On the other hand, after the phase transition, the inter-module connections facilitate global and internal synchrony, as they bring modules close together. Figures 2(a) and 4(b)-(c) show that the parameter values at which the system transitions is the same as the equivalent $s$ oscillators system, validating the coarse-graining method. Nonetheless, assuming high modular synchrony as we did is a strong assumption because it is highly dependent upon the number of connections between modules, as we see from Figure 2 (b), where fewer connections creates a demand for higher $\lambda_{\text{in}}$. We naturally expect this structural role to also be contained in inner-modular connections, with more connections within modules positively correlating to lower internal coupling.
We also applied the method to real networks where modules are not so well defined. We considered a small network with two modules, Zachary’s Karate club social network, and the C. Elegans gap junctions neural network, which can be modularized in different ways. Our results, displayed in Figure 5, show that the complexity these networks introduces difficulties that cannot be captured by the simple coarse-graining processes proposed here. The reduction of whole modules into single oscillators is not sufficient to describe the synchronization process in detail. However, some intuition can still be gained by the method, such as the approximate value of critical coupling and even the behavior of the order parameter if the internal coupling is large enough. Yet, the inaccuracy of the approximation raises the question of how reliable are the dynamics inferred from networks obtained from measurements of brain activity, e.g., where large areas of the brain, not necessarily comprising a module, are replaced by a single node, such as in functional magnetic resonance imaging, near-infrared spectroscopy and electroencephalograms.

### v2

@@ IV Two modules
Figure 2: Critical inter-module (a) and inner-module (b) coupling strengths for fully connected modules as a function of coupling probabilities $p=0.01;0.25;0.5;0.75;1$ and frequencies $\omega=1.5;2.5;3.5$. Lines represent the corresponding 2 oscillator system. In all simulations $N_{1}=200$ and $N_{2}=100$; Gaussian frequency distributions with $\Delta=1$ and mean $(\omega_{1},\omega_{2})=(\omega,0)$. Error bars made over 3 simulations.
The results are very similar to the fully connected case, although requiring larger values of $\lambda_{\text{in}}$ to reach the same internal synchronization threshold of $q_{\sigma}=0.9$. The coarse-graining procedure is, therefore, largely independent of the network structure, as expected from the theory. However, we note that the validity of the process relies on the synchronizability of the modules, which is not always attainable. Ling et al. showed that for an Erdös-Rényi network the connectivity threshold for synchronization is $p\sim\log{(n)}/n$ which is about 0.03 for $n=200$ [49, 50]. Even though our simulated Erdös-Rényi network has different couplings to account for modularity, we confirmed that for small enough connection probability ($p=0.01$) modules do not synchronize regardless of $\lambda_{\text{in}}$ (simulations not shown). In section VI we will encounter this situation again.
The onset of synchronization indicates both minimum $\lambda_{\text{in}}$ for $r_{\sigma}\geq q_{\sigma}$ and $\lambda$ for global synchrony. Figure 2 shows such $(\lambda_{\text{in}},\lambda)$ pairs for 5 different values of frequency offsets $\omega$ and 5 different inter-module connection probabilities $p$ for fully connected modules. Panel (a) shows that, for large enough $\lambda_{\text{in}}$, the global phase transition happens at about $\lambda=\omega$ which is precisely when the phase transition of a 2 oscillator system takes place, in accordance with the coarse-graining process. More interesting, the transition point does not depend upon the number of connections between modules, as expected from Eq. (6). Panel (b) shows that immediately after the transition to the globally synchronized state, the minimum $\lambda_{\text{in}}$ that would assure sufficient synchrony for the coarse-graining process decreases non linearly with the amount of connections between the modules, exposing the importance of network structure in keeping inner-module cohesion.
V Three modules
We now extend the results of the previous section to systems with three modules, as there are still some analytical tools available to represent the states of three Kuramoto oscillators. We divide the work into two parts. First, we find an approximation for the global phase transition in parameter space, Eq. (22). Secondly, we write Eq. (11) for 3 modules to find upper and lower bounds within the synchronization region.
V.1 Three oscillators
For 3 oscillators the Kuramoto model can be written in terms of the phase differences $\phi_{12}\equiv\theta_{1}-\theta_{2}$ and $\phi_{23}\equiv\theta_{2}-\theta_{3}$ as
@@ V.1 Three oscillators
Similar to two oscillators, the order parameter can be written in terms of phase differences as
@@
Figure 3 (a) shows the contour plot of Eq. (18) in the phase space $\phi_{12}\times\phi_{23}$. To determine the stability of these fixed points we calculated the Jacobian of Eqs. (16) and looked at the largest real part of the pair of eigenvalues: when it is zero we are at the critical line, shown as the red curve in Fig.3 (a) . The curve divides the phase space into two regions: inside where the system displays stable synchrony, and outside, where motion is periodic. More elucidating, however, is to understand synchronization in the parameter space $\Omega_{1}\times\Omega_{3}$. Figure 3 (b) shows the stable $r$ values from panel (a) reparameterized according to Eq. (17). It is interesting to note that, regardless of being displayed in the phase space or the parameter space, contours of constant synchronization resemble ellipses. We proceed to finding approximations for such contours.
Figure 3: (a) Contour plot of Eq. (18) in the phase space $\phi_{12}\times\phi_{23}$. The red curve divides stable (inside) from unstable (outside) points. (b) Stable points in the $\Omega_{1}\times\Omega_{3}$ plane according to Eq. (17). (c) $r$ values along the ellipses. Dashed lines are the empirical values from Eq. (21). (d) Contrast between the ellipses (Eq.(20), dashed) and lines of constant $r$ (continuous).
@@
In the inset of Figure 3 (d) we can easily see that there are constant $r$ values that fall out of the theoretical ellipses, therefore, there should be a smaller ellipse that encompass all values. We numerically find that further increasing $\alpha$ up to $\alpha_{2}\equiv 9.295$ causes all possible synchronization outcomes to tightly fall within the ellipse. In other words, depending on how we control our parameters $\{\Omega_{1},\Omega_{2}\}$ the phase transition happens somewhere between $\alpha_{1}\leq\alpha\leq\alpha_{2}$, or in terms of the original parameters,
$\lambda^{2}\leq\omega_{1}^{2}-\omega_{1}\omega_{3}+\omega_{3}^{2}\leq 1.0328\lambda^{2}.$ (22)
This is an estimate for the phase transition of 3 oscillators. As we argued in previous sections, both the coarse-grained system and the equivalent three oscillators system have the same phase transition due to the form of Eq. (6). However, it is important to emphasize that so far we do not have an approximation for $r$ within the ellipse that takes into account modules sizes and synchronization thresholds.
V.2 Coarse graining for three modules
Writing Eq. (11) for 3 modules we obtain
$r^{2}=\rho_{1}^{2}r_{1}^{2}+\rho_{2}^{2}r_{2}^{2}+\rho_{3}^{2}r_{3}^{2}+2\rho_{1}\rho_{2}r_{1}r_{2}\cos{(\phi_{12})}+2\rho_{1}\rho_{3}r_{1}r_{3}\cos{(\phi_{12}+\phi_{23})}+2\rho_{2}\rho_{3}r_{2}r_{3}\cos{(\phi_{23})},$ (23)
from which we can write $r_{+}$ and $r_{-}$. Despite having an expression for the order parameter in terms of the stable phase differences we do not have explicit solutions for such phases, only implicit ones given by the set (17). Consequently, determining a theoretical closed expressions as we did in Eq. (15) is not possible. Our approach for 3 modules thus focuses on writing $r_{+}$ and $r_{-}$ via numerically solving system (17) to substitute into Eq. (23). Note that the particular case of $\rho_{i}=1/3$ and $r_{i}=1$ throws us back to the 3 oscillators order parameter, Eq. (18).
@@ V.2 Coarse graining for three modules
Figure 4: Simulation of Eq. (4) with 3 modules. Panels (a)-(c) show final values of $\lambda_{\text{in}}$, global order parameter $r$ and standard deviation $\sigma_{r}$ in the frequency parameter space. Panels (d) and (e) show numerical estimates of $r_{+}$ and $r_{-}$ taken from Eq. (23) with $r_{\sigma}=1$ and $r_{\sigma}=q_{\sigma}=0.9$, respectively. Phase differences were obtained numerically, solving system (17). In all panels the black ellipse is the lower limit of inequality (22) for the phase transition. Simulation values: $N_{1}=90$, $N_{2}=100$, $N_{3}=110$, $p=0.01$ and $\lambda=3$. Natural frequencies are Gaussian distributed with $\Delta=1$ and mean $\omega_{1}$, $\omega_{2}=0$ and $\omega_{3}$.
@@
We now proceed to simulate Eq. (4) with three modules and, as before, we take modules of distinct sizes to be more realistic, namely $N_{1}=90$, $N_{2}=100$ and $N_{3}=110$. Throughout this section we will work with fully connected modules, $p=0.01$ and $\lambda=3$ so that $\Omega_{i}=\omega_{i}$. Again, we set the frequency distributions to be Gaussians of width $\Delta=1$ centered at $\omega_{1}$, $\omega_{2}=0$ and $\omega_{3}$. The minimum synchrony of each module is fixed at $q_{\sigma}=0.9$.
@@
Figure 4 (a) shows the final value of $\lambda_{\text{in}}$. We see the same behavior observed in the case of two modules: approaching the synchronization regime from a asynchronous state increases the internal coupling required to preserve $r_{\sigma}\geq q_{\sigma}$. This might seem an odd, considering we are decreasing the absolute frequency value and consequently expect it to be easier to keep cohesive modules. However, given that there are few connections between modules, the transfer of information from one to the others is harder when their frequencies are further apart, even though the absolute value is smaller.
Figures 4 (b)-(c) show the global order parameter and its standard deviation, respectively. The black curve shows the lower elliptic bound given by Eq.(22). We observe that the synchronization region corresponds to the expected behavior for three oscillators and exhibit all distinct behaviors, full and partial phase locking and asynchrony. Panels (d) and (e) show the numerical estimates of the synchronization bounds $r_{+}$ and $r_{-}$ taken from Eq. (23), respectively. The phase differences were obtained by numerically solving system (17). Note that the simulated value of $r$ appears to fall in between both, as expected.
VI Real Networks
@@
We have seen so far that, given high enough $\lambda_{\text{in}}$, the coarse-graining procedure recreates the synchronization properties of the reduced system in artificial modular networks with fully connected and Erdös-Rényi modules (shown only for the case of two modules). These networks, however, were generated in such a way as to exhibit very strong modular structure that does not resemble real world systems. In what follows we test our results on two well known networks, Zachary’s Karate club social network [51] and the C. Elegans gap junctions neural network [26]. Both of these networks can be modularized using standard procedures. What distinguishes them from the artificial networks constructed in this paper are structural properties such as not having full number of connections within modules and non random connections between modules per se.
@@
Figure 5: Dynamics of Eq. (4) for Zachary’s Karate social network. (a): Final $\lambda_{\text{in}}$ and a function of $\lambda$. Magenta and turquoise lines marks the average of critical $\lambda$ and $\lambda_{\text{in}}$. (b): $\langle r\rangle$ with $\lambda_{\text{in}}=7.9$ (the critical value) in black and with $\lambda_{\text{in}}=20$ in red. Green solid curves represent the theoretical results, Eqs. (12) and (13). Simulation values: $\Delta=0$, $\omega_{1}=1.5$ and $\omega_{2}=0.0$.
@@
We start with the Karate network [51], that has a total of 34 nodes with 78 connections. To modularize the network we used Girvan-Newman’s method, splitting it in two modules of 17 nodes each [52]. Connections are divided into 32 and 35 for each module and 11 connecting the modules. As before, we run the dynamics given by Eq. (4) on this network and slowly increase $\lambda_{\text{in}}$ until a minimum inner module synchrony of 90% is reached. Figure 5 (a) shows the minimum internal coupling that keeps the modular synchrony above the threshold. Note that the global phase transition, expected at $\lambda=1.5$ given the strong bi-modular characteristic of Zachary’s Karate network, ocurs at $\lambda=1.65$. Just as in the artificial network, increasing $\lambda$ for asynchronized states impairs modular synchrony, hence the continuous increase in $\lambda_{\text{in}}$ as $\lambda$ grows. On the other hand, after the phase transition, synchronized states act in cohesion, requiring lesser $\lambda_{\text{in}}$ as we increase $\lambda$. However, the slopes of increasing and decreasing values of $\lambda_{\text{in}}$ are quite different from what we found in section IV. This behavior reflects the quantitative differences between $\lambda_{\text{in}}$ values among the real and artificial networks, Figures 1(c) and 5(a). The information transfer among modules is harder in Karate’s network, which has proportionally fewer connections among modules, 11 out of 289, corresponding to $p\approx 0.04$, as opposed to $p=0.1$ for the artificial fully connected network. Consequently, Zachary’s Karate network requires much lower $\lambda_{\text{in}}$ prior to the phase transition and higher after it.
The order parameter as a function of $\lambda$ is shown in Figure 5(b). We fixed the internal coupling to the minimum value that assures modular synchrony, i.e., $\lambda_{\text{in}}=7.9$, corresponding to the peak in panel (a). The order parameter (black points), shows that synchronized states do not reach the $r_{+},r_{-}]$ region, in green. This is likely due to the low connectivity between modules which lead to large inter-modular phase difference. The global phase transition point is also slightly larger than the expected value of 1.5. We can counteract this structure effect by overshooting the internal coupling. The red curve of Figure 5 (b) shows the synchronization values when keeping $\lambda_{\text{in}}=20$ throughout the whole simulation. Forcing high inner-module coupling strength pushes both modules to highly synchronized states so that the system behaves almost exactly as two oscillators.
As for our second example we study a reduced network extracted from the C. Elegans gap junctions neural network containing 248 nodes and 511 connections [53]. Using appropriate metrics it can be modularized into 3, 5 or 10 modules [54, 55, 56]. Note that regardless of number of modules, some end up being far greater than others, for instance, in the 3 modules arrangement we have 130 nodes in the largest module, 77 nodes in the intermediate and 41 nodes in the smallest. For 5 modules the number of nodes in each module are 120, 83, 34, 7 and 4 and with 10 modules the module sizes are 68, 33, 76, 11, 13, 7, 6, 8, 8 and 18. The modularity coefficient [52] for these partitions is $Q=0.44$ for 3 and 5 modules and $Q=0.55$ for 10 modules [55]. Moreover, the network has about 1% of the total number of possible connections, indicating that it is hard to make all modules have a minimum synchrony of 90% or even lower. For that reason we made some simplifying assumptions to test if the coarse graining procedure could work at least under favorable conditions. First, we assumed that oscillators in each module are identical, i.e. $\Delta=0$. Second, given the small density of connections in some of the modules, we attributed initial phases for the oscillators in the interval $0<\theta_{\sigma,i}<\epsilon$, where $\epsilon$ was typically around $\pi$. This condition guarantees that the initial conditions are within the basin of attraction of the synchronized state for each module. That, however, does not imply that all modules will reach the desired synchronization level, as they are coupled to the rest of the network. The values of natural frequencies for the modules where chosen in a rather arbitrary way (see caption of Figure 6 for specific values) to produce a non-trivial global order parameter.
Figure 6: Dynamics of Eq. (4) for C. Elegans gap junctions network. (a), (b) and (c) show $\langle r\rangle$ for networks clustered into 3, 5 and 10 modules with $\lambda_{\text{in}}=20$ (red) and coarse grained prediction $r_{-}$ (green). Simulation values: $\Delta=0$, (d): $(\omega_{1},\omega_{2},\omega_{3})=(0,1,6)$; (e): $(\omega_{1},\omega_{2},\omega_{3},\omega_{4},\omega_{5})=(0,1,1,-2,7)$; (f): $(\omega_{1},\omega_{2},\omega_{3},\omega_{4},\omega_{5},\omega_{6},\omega_{7},\omega_{8},\omega_{9},\omega_{10})=(0,1,-3.5,-2,2,0.6,0.4,1.2,1.5,-1.5)$. panels (d) and (e) show the coupled ($\lambda=2$, $\lambda_{\text{in}}=20$) and decoupled ($\lambda=0$, $\lambda_{\text{in}}=1$) dynamics of 3 modules for same initial phases. Panel (f) shows the maximum angular spread of initial phases $\epsilon$ such that the decoupled modules synchronize above 90%.
Before we compare exact and coarse-grained results for the C. Elegans we emphasize that, although the network’s backbone is always the same (same nodes and same connections) each partition of the network into modules leads to a different distribution of interaction strengths $\lambda_{\text{in}}$ (within modules) and $\lambda$ (between modules). Therefore each case responds to the Kuramoto dynamics differently.
Figure 6(a)-(c) show in red the simulation of $\langle r\rangle$ for the 3 modularizations above, each accompanied by a simulation of the theoretical order parameter $r_{-}$ of the reduced system in green for comparison. For $r_{-}$ the phase differences were obtained by integrating the respective $s$ oscillators system and using for $q_{\sigma}$’s the mean values of $r_{\sigma}$ obtained from the full network simulation as in Fig. 6(d). In all cases the agreement is surprisingly good, despite the complex behavior and the relatively low value of global synchronization displayed by $\langle r\rangle$. Panels (d)-(f) show the details of the method for the case of 3 modules. In (d) we see how module 1 (red) reaches $q_{1}\approx 0.8$, module 2, $q_{2}\approx 0.7$, and module 3 does not synchronize even for large $\lambda_{\text{in}}$, attaining $q_{3}\approx 0.2$ only. The green points in (a) were obtained using these values for the $q_{\sigma}$’s in Eq. (13). Panel (e) shows the order parameters for the modules for $\lambda=0$ using the same initial conditions as in (d), indicating that they are synchronizable when independent, but that the interaction messes with their internal synchrony. Finally, panel (f) shows, also for $\lambda=0$, how the maximum spread $\epsilon$ of initial phases that allow convergence to the synchronized state increases with $\lambda_{\text{in}}$. We see, in particular, that module 3 only synchronizes if $\epsilon<2.5$, even for identical oscillators and $\lambda=0$, indicating that the synchronized state is stable but with a not so large basin of attraction. A similar study was performed for the 5 and 10 modules partitions of the network, the corresponding $q_{\sigma}$’s were used to compute $r_{-}$.
VII Discussion
In this work we proposed a coarse-graining method for modular networks where each module is replaced by a single effective node, drastically reducing the network size. The order parameter of the reduced network takes into account the relative size of the module represented by the effective node and its average degree of internal synchronization. To test the coarse-graining procedure we initially constructed artificial networks under two core assumptions: first, modules should be synchronizable, in the sense that for sufficiently large internal coupling $\lambda_{\text{in}}$ the local order parameter for the module should reach values $r_{\sigma}$ of the order of $q_{\sigma}=0.9$. This ensures that the hypothesis of phase synchronization within the module is approximately true. Second, all intra-module connections are fixed to the same value $\lambda_{\text{in}}$, while inter-module connections take a value $\lambda$, which we vary to observe the transition to global synchronization. For synthetic networks with 2 and 3 modules the procedure works very well, as indicated in Figures 1 to 4. We tested the approach for fully connected and randomly connected modules.
For the two real networks considered in section VI we relaxed the first condition, allowing $r_{\sigma}$ to be smaller than the desired value $q_{\sigma}=0.9$, as small modulus tend not to synchronize even for large internal coupling strength. That, however did not spoil the coarse graining, most likely because small modulus give small contributions to $r_{\pm}$. Taking this into account we showed that results for the C. Elegans were surprisingly good for all partitions considered. We presume that this excellent agreement was possible because we considered identical oscillators in each module and chose initial phases not too far away from synchrony, facilitating synchronization for most of the modules. Further studies considering distributions of natural frequencies (with different average values) and random initial phases (as we did in sections IV and V for artificial networks) would be important to understand the robustness of our results.
Given the experiments with artificial and real networks, we can summarize the coarse graining method according to the following recipe: (i) for a given network with $s$ modules, fix a desired synchronization threshold $q_{\sigma}$; (ii) for each inter-module coupling $\lambda$, run the Kuramoto model with Eq. (4) and vary the internal coupling $\lambda_{\text{in}}$ until $r_{\sigma}\geq q_{\sigma}$; (iii) from $\lambda_{\text{in}}$ as a function of $\lambda$ take the largest $\lambda_{\text{in}}$, which should correspond to $\lambda$ close to the phase transition of the network; (iv) if one or more modules do not reach the desired synchronization threshold $q_{\sigma}$, replace $q_{\sigma}$ by $\langle r_{\sigma}\rangle$ for these modules; (v) run the standard Kuramoto dynamics on the reduced network with $s$ nodes and compute the asymptotic phase differences $\phi_{\sigma\sigma^{\prime}}$; (vi) compute $r_{\pm}$ using Eqs. (12) e (13) and compare with the order parameter of the original network.
It would also be interesting to compare our results with the Ott-Antonsen theory (OA) [57] applied to modular networks. The OA ansatz reduces the dynamics of $s$ modules to $2s$ coupled differential equations, for module and phase of the order parameters computed for each module separately, from which the global order parameter can be calculated. [58]. However it is important to keep in mind that such procedure works only for infinitely many oscillators and Lorentzian distribution of natural frequencies. Our approach differs from OA in the sense that we replace modules of the original network by effective oscillators and run the Kuramoto model on this smaller system. Although the method was inspired by EEG measurements, it does not require infinitely many oscillators per module nor specific distributions of natural frequencies. However, we do required modules to be ”synchronizable”.
We should mention that we require prior knowledge of the network structure. In particular, we assume the network to be static in its composition and, therefore, the method does not accommodate merging or splitting of modules. Also, the fact that $\lambda_{\text{in}}$ is generally much larger than $\lambda$ in the vicinity of the phase transition ensures that inter modular dynamics evolve on a faster time scale than that of the full network, justifying the replacement of modules by effective nodes.